Year ZAR/GBP
1961-1982 2
1982-1985 1.95
1985 2.99
1986-1988 3.47
1989-1993 4.51
1994 5.41
1999 9.07
2001 21.3
2002 13.85
2005 10.2
2008 17.82
2010 12.01
2015 18.27

What is the rate of depreciation from 1961 till 1994?

Please explain the formula used and what it is called.
Answer #1

In 1961: we needed 2 ZAR to exchange 1 GBP. In 1994 we needed 5.41 ZAR to exchange 1 GBP. This means GBP has become expensive and ZAR has deteriorated or depreciated.

Let's invert the rate. In 1961, we needed 2 ZAR to get 1 GBP. This implies we got 0.5 GBP for 1 ZAR. In 1994, we needed 5.41 ZAR to exchange 1 GBP. This means, in 1 ZAR, we can buy only 1/5.41 = 0.1848 GBP

Hence, depreciation = (0.5 - 0.1848) / 0.5 = 63.03%

And annual rate = 63.03% / (1994 - 1961) = 1.91%